OpenStudy (anonymous):

estemate 1/8 + 3/5 a. 1/2 b. 1/3 c. 1/4 d. 2/3

OpenStudy (anonymous):

d

OpenStudy (anonymous):

cause the exact result is 29/40

OpenStudy (anonymous):

sanks!

OpenStudy (anonymous):

denada

OpenStudy (anonymous):

\[\frac{a}{b} + \frac{c}{d} = \frac{ad+ cb}{bd}\]